# From Cool Demos to Production-Ready FMware: Core Challenges and a Technology Roadmap

### v1

@@ Abstract
The rapid expansion of foundation models (FMs), such as large language models (LLMs), has given rise to FMware—software systems that integrate FMs as core components. While building demonstration-level FMware is relatively straightforward, transitioning to production-ready systems presents numerous challenges, including reliability, high implementation costs, scalability, and compliance with privacy regulations. This paper provides a thematic analysis of the key obstacles in productionizing FMware, synthesized from industry experience and diverse data sources, including hands-on involvement in the Open Platform for Enterprise AI (OPEA) and FMware lifecycle engineering. We identify critical issues in FM selection, data and model alignment, prompt engineering, agent orchestration, system testing, and deployment, alongside cross-cutting concerns such as memory management, observability, and feedback integration. We discuss needed technologies and strategies to address these challenges and offer guidance on how to enable the transition from demonstration systems to scalable, production-ready FMware solutions. Our findings underscore the importance of continued research and multi-industry collaboration to advance the development of production-ready FMware.
I Introduction
@@
FMware refers to software integrating foundation models (FMs), like Large Language Models (LLMs), as core components [1]. Since ChatGPT’s release in late 2022, the FMware landscape has exploded, with over 600,000 open-source models, including FMs and other AI models, now available on platforms like Hugging Face [2]. Goldman Sachs predicts FMware could boost global GDP by 7% [3].
@@
Building impressive demos with FMs is relatively easy, but transitioning to production-ready FMware incurs significant challenges due to its inherent complexity as a compound system [4, 1]. Unlike static systems, FMware integrates multiple dynamic components for real-time updates, control, and adaptability [4, 5]. This complexity is reflected in industry adoption rates: a survey of 430 technology professionals revealed that only 10% of organizations had launched FMware in production environments [6]. Respondents identified reliability, high implementation costs, latency, compliance, and privacy as the top challenges preventing them from moving from demos to production-ready FMware.
@@
The transition to production-ready FMware is formidable due to its dynamic nature, which combines classical software engineering complexities with unique issues inherent to foundation models. Production-ready FMware must continuously evolve to meet customer expectations, requiring consistent performance, reliability, feature updates, and adherence to service level agreements (SLAs), all while managing operational costs. For instance, OpenAI’s infrastructure costs for running ChatGPT in 2023 were estimated at \$700,000 per day [7]. Beyond these traditional challenges, FMware developers also face unique problems like handling hallucinations, increased inference costs, and orchestrating tasks across various AI components [8, 1]. LinkedIn’s experience illustrates the effort required to reach production readiness: they achieved 80% functionality in a month but spent four more months to complete the remaining 20%, with diminishing returns on each additional 1% improvement [9]. Similarly, Microsoft and GitHub found that testing FMware becomes prohibitively expensive as complexity scales [10]. These real-world examples underscore the need for robust, system-based approaches to build reliable, production-ready FMware solutions [4, 8, 1].
@@
In this paper, we employ a thematic analysis to systematically identify key challenges in developing production-ready FMware. We draw on insights from multi-industry collaborations, conferences, customer meetings, hands-on development of the FMware lifecycle platform (FMArts) [1] and literature surveys. In doing so, we provide the first comprehensive articulation of the recurrent issues across different stages of FMware development, consolidating them into challenges that hinder productionizing FMware.
@@
This paper complements our vision for Software Engineering 3.0 (SE 3.0), where we advocate for an AI-native, intent-first approach where AI systems evolve from task-driven copilots to intelligent collaborators [11]. Both SE 3.0 and our paper address the need to adapt software engineering practices to manage the complexity and dynamism of AI systems like FMware. By synthesizing hands-on knowledge from building production-ready FMware and interacting with industry experts, we underscore the urgency of addressing these challenges to move from “cool” demos to robust, production-ready solutions. Our study further highlights the importance of ongoing research and multi-industry collaboration as FMware continues to evolve.
@@
This paper is organized as follows: Section II covers the evolution of FMware and related work. Section III explains our thematic analysis approach. Section IV outlines the FMware lifecycle and spot recurring issues at each stage that affect production readiness. Section V presents the key challenges hindering FMware production-readiness and suggests future directions. Finally, Section VI concludes the study.
II Related Work
@@
Despite a wealth of research on software release readiness [12, 13, 14, 15, 16], there is a notable lack of studies on production readiness [17, 18]. Release readiness ensures that software meets functional requirements and is ready for distribution. Whereas, production readiness also guarantees reliable operation in live environments by addressing factors like latency and operational costs. Early work by Mockus [19] explored an analogy-based prediction method to model and forecast the workflow of work items in software projects to asses release readinees of a software. Port and Wilf [15] explored release readiness through a case study at NASA JPL, followed by several studies quantifying and predicting release readiness factors using machine learning [12, 13, 14]. More recently, Patel et al. [16] proposed a checklist to assess the release readiness of FMware.
@@
In assessing production readiness, Asthana and Olivieri [17] developed a metric focused on software quality and reliability, while Cusick [18] proposed an architecture review method. More recently, Breck et al. [20] in their seminal work, examined the production readiness of machine learning systems, proposing a rubric to reduce technical debt and evaluate readiness. Parnin et al. [10] documented the difficulties of building a production-ready copilots, by interviewing the team that built various copilots at Microsoft. In their influential blog post, Yan et al. [21] shed light on some of the challenges that impact the production-readiness of FMware. However, to the best of our knowledge, our paper is the first work to conduct a systematic study to identify the associated challenges with building production-ready FMware.
III Methodology
We used thematic analysis on various sources of data, including conference reports, meeting notes, and collaborative discussions to identify key challenges in developing production-ready FMware. While traditionally applied in software engineering studies [22, 23] to analyze structured or semi-structured interviews, thematic analysis is also well-suited for extracting themes from unstructured data, such as that we utilized [24]. Thematic analysis typically involves coding the data, identifying patterns, and grouping codes into overarching themes, which allows us to follow a structured approach to ensure the systematic identification of challenges.
@@ III Methodology
Step 1: Data Collection. We collected data from various sources to gain insights into the challenges that researchers and practitioners encounter when productionizing FMware.
@@
  • OPEA Project Participation: As active participants, we played a significant role in the OPEA initiative (Open Platform for Enterprise AI (OPEA) [8]). In particular, the last author leads OPEA’s research working group. OPEA is a global multi-company collaborative initiative focused on addressing the challenges of making FMware enterprise-ready. OPEA provides detailed frameworks, architectural blueprints, and a four-step assessment for evaluating FMware in terms of enterprise readiness. We used meeting notes (some of which are publicly available [25]) from OPEA discussions, which involved 43 organizations (e.g., Intel, AMD, RedHat, Docket, JFrog, Anyscale, LlamaIndex and SAP).
@@
  • SPDX Working Group: First author co-led SPDX community meetings (an ISO/IEC 5692:2021 Software Bill of Materials standard [26]) over the past three years, contributing to discussions on the AI Bill of Materials (AI BOM), which included discussing challenges in building production-ready FMware. The meeting notes, available publicly [27], were incorporated into our analysis.
@@
  • Inception and Launch of the AIware Conferences, summit and Bootcamp: Authors of this paper spearheaded the first ACM International Conference on AI-Powered Software (AIware) [28], co-located with FSE 2024, the FM+SE summit series [29] and the AIware Bootcamp [30]. Engaging with participants from Google, Microsoft, GitHub, and other industry and academic professionals provided valuable insights into the challenges of productionizing FMware, which we summarized in a report used in our analysis [31].
@@
  • Internal Industry Experience: Our collaborations with customers and development teams to understand FMware’s functional and non-functional requirements. Additionally, all authors contributed to developing and productionizing FMArts, a comprehensive lifecycle engineering platform for FMware [1]. Our practical experience in creating FMware with FMArts and making it production-ready [1] and integrating it into Huawei products (e.g., Huawei Cloud) informed the thematic analysis.
  • In-depth Literature Review: We began with a review of grey literature, including blog posts, whitepapers, and developer forums, and noticed recurring discussions about the production readiness of FMware. Building on these insights, we conducted a focused review of academic papers to find supporting evidence. Through this activity, we provide a comprehensive analysis of the challenges in deploying FMware, grounding all issues in Section IV with citations and minimizing subjective bias.
Step 2: Data Coding and Initial Thematic Grouping. After data collection, three of the four authors collaborated to conduct the initial coding. Each dataset (conference reports, meeting notes, and collected grey literature) was systematically reviewed, and recurrent issues related to different stages of FMware lifecycle (Figure 1) were coded. An example code for an identified issue in the Data and FM Alignment stage would be low data quality. We took special care to not include the issues that were common to both demo and production-ready FMware.
Step 3: Collaborative Discussion to Identify Overarching Themes. In the next step, three of the four authors collaborated to refine the codes and identify overarching themes (which we present as challenges). Similar codes were grouped into higher-order themes through iterative discussions. These themes spanned multiple FMware lifecycle stages, presenting a holistic view of the involved challenges in moving from demo systems to production-ready FMware.
Step 4: Thematic Consolidation. Finally, we revisited and refined the overarching themes to ensure they were representative of the data and accurately captured the challenges. We grouped the issues into clear, concise themes, which are presented in Section V. This step allowed us to provide a clear, systematic view of the critical challenges in making FMware production-ready.
IV Recurrent Issues in Productionizing FMware
@@
Figure 1: FMware Lifecycle.
IV-A FM(s) Selection
The first step in developing FMware is to select the FM(s) to be used. There exist over 600,000 open-source models available (among FMs and the more traditional machine learning models) [2]. Typically, the number of FMs to be selected depends on the complexity and functionality required by the FMware. For example, a simple chatbot may only need one FM, while complex systems might require multiple FMs for its different functionalities.
Difficulty balancing functional requirements with performance and costs. Selecting FMs for FMware requires balancing functionality with performance, infrastructure, and costs [32]. Demos favor large general-purpose models like ChatGPT-4, neglecting latency and resource limits due to the typically small number of required API calls. LLaMa-70B, for example, needs 140GB VRAM to be hosted locally [33]. In production, smaller FMs are preferred for cost, latency, and privacy reasons (especially in the context of edge devices, such as smartphones) [34, 35], though they may lack the full capabilities of larger FMs [36, 37]. However, it is important to note that even hosting smaller models can also increase ownership costs [32].
IV-B Data and FM Alignment
After selecting FMs, the next crucial step is data and model alignment. This process tailors the FM(s) to meet the specific requirements and goals of a given FMware application [38].
Data alignment curates specialized data through labeling, generation, debugging, and testing, especially for enterprise cases [39, 40]. Domain-specific datasets are created to ensure that an FMware can perform well on the use-case specific data. Model alignment adjusts FM parameters using techniques like supervised fine-tuning and Reinforcement Learning with Human Feedback (RLHF) to ensure the FMware behaves as intended [41].
Low data quality. High data quality is crucial for reliable FMware, as poor data leads to untrustworthy outputs [40].

### v2

@@ Abstract.
The rapid expansion of foundation models (FMs), such as large language models (LLMs), has given rise to FMware—software systems that integrate FMs as core components. While building demonstration-level FMware is relatively straightforward, transitioning to production-ready systems presents numerous challenges, including reliability, high implementation costs, scalability, and compliance with privacy regulations. Our paper conducts a semi-structured thematic synthesis to identify the key challenges in productionizing FMware across diverse data sources including our own industry experience in developing FMArts—a FMware lifecycle engineering platform and integrating it into Huawei cloud, grey literature, academic publications, hands-on involvement in the Open Platform for Enterprise AI (OPEA), organizing the AIware conference and Bootcamp, and co-leading the ISO SPDX SBOM working group on AI and datasets. We identify critical issues in FM selection, data and model alignment, prompt engineering, agent orchestration, system testing, and deployment, alongside cross-cutting concerns such as memory management, observability, and feedback integration. We discuss needed technologies and strategies to address these challenges and offer guidance on how to enable the transition from demonstration systems to scalable, production-ready FMware solutions. Our findings underscore the importance of continued research and multi-industry collaboration to advance the development of production-ready FMware.
AI-powered Software, Production-ready FMware, FMware, Productionization
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligenceccs: Software and its engineering
1. Introduction
@@
FMware refers to software integrating foundation models (FMs), like Large Language Models (LLMs), as core components (Hassan et al., 2024b). Since ChatGPT’s release in late 2022, the FMware landscape has exploded, with over 600,000 open-source models, including FMs and other AI models, now available on platforms like Hugging Face (Santa, [n. d.]). Goldman Sachs predicts FMware could boost global GDP by 7% (gol, [n. d.]).
@@
Building impressive demos with FMs is relatively easy, but transitioning to production-ready FMware incurs significant challenges due to its inherent complexity as a compound system (Zaharia et al., 2024; Hassan et al., 2024b). Unlike static systems, FMware integrates multiple dynamic components for real-time updates, control, and adaptability (Zaharia et al., 2024; ibm, [n. d.]). This complexity is reflected in industry adoption rates: a survey of 430 technology professionals revealed that only 10% of organizations had launched FMware in production environments (int, 2023). Respondents identified reliability, high implementation costs, latency, compliance, and privacy as the top challenges preventing them from moving from demos to production-ready FMware.
@@
The transition to production-ready FMware is formidable due to its dynamic nature, which combines classical software engineering complexities with unique issues inherent to foundation models. Production-ready FMware must continuously evolve to meet customer expectations, requiring consistent performance, reliability, feature updates, and adherence to service level agreements (SLAs), all while managing operational costs. For instance, OpenAI’s infrastructure costs for running ChatGPT in 2023 were estimated at \$700,000 per day (Abbas, 2024). Beyond these traditional challenges, FMware developers also face unique problems like handling hallucinations, increased inference costs, and orchestrating tasks across various AI components (OPEA, [n. d.]; Hassan et al., 2024b). LinkedIn’s experience illustrates the effort required to reach production readiness: they achieved 80% functionality in a month but spent four more months to complete the remaining 20%, with diminishing returns on each additional 1% improvement (lin, 2024). Similarly, Microsoft and GitHub found that testing FMware becomes prohibitively expensive as complexity scales (Parnin et al., 2023). These real-world examples underscore the need for robust, system-based approaches to build reliable, production-ready FMware solutions (Zaharia et al., 2024; OPEA, [n. d.]; Hassan et al., 2024b).
@@
In this paper, we employ a semi-structured thematic synthesis (more details in Section 3) to systematically identify key challenges in developing production-ready FMware. We draw on insights from multi-industry collaborations, conferences, customer meetings, hands-on development of the FMware lifecycle platform (FMArts) (Hassan et al., 2024b) and literature surveys. In doing so, we provide the first comprehensive articulation of the recurrent issues across different stages of FMware development, consolidating them into challenges that hinder productionizing FMware.
@@
This paper complements our vision for Software Engineering 3.0 (SE 3.0), where we advocate for an AI-native, intent-first approach where AI systems evolve from task-driven copilots to intelligent collaborators (Hassan et al., 2024c). Both SE 3.0 and our paper address the need to adapt software engineering practices to manage the complexity and dynamism of AI systems like FMware. By synthesizing hands-on knowledge from building production-ready FMware and interacting with industry experts, we underscore the urgency of addressing these challenges to move from “cool” demos to robust, production-ready solutions. Our study further highlights the importance of ongoing research and multi-industry collaboration as FMware continues to evolve.
@@
This paper is organized as follows: Section 2 covers the evolution of FMware and related work. Section 3 explains our thematic analysis approach. Section 4 outlines the FMware lifecycle and spot recurring issues at each stage that affect production readiness. Section 5 presents the key challenges hindering FMware production-readiness and suggests future directions. Finally, Section 6 concludes the study.
2. Related Work
2.1. Release and Production Readiness in Software Engineering
@@
Despite a wealth of research on software release readiness (Al Alam et al., 2014; Koprowski et al., 2014; Al Alam et al., 2017; Port and Wilf, 2013; Patel et al., 2024), there is a notable lack of studies on production readiness (Asthana and Olivieri, 2009; Cusick, 2013). Release readiness ensures that software meets functional requirements and is ready for distribution. Whereas, production readiness also guarantees reliable operation in live environments by addressing factors like latency and operational costs. Early work by Mockus (2003) explored an analogy-based prediction method to model and forecast the workflow of work items in software projects to asses release readinees of a software. Port and Wilf (2013) explored release readiness through a case study at NASA JPL, followed by several studies quantifying and predicting release readiness factors using machine learning (Al Alam et al., 2014; Koprowski et al., 2014; Al Alam et al., 2017). More recently, Patel et al. (2024) proposed a checklist to assess the release readiness of FMware.
@@
In assessing production readiness, Asthana and Olivieri (2009) developed a metric focused on software quality and reliability, while Cusick (2013) proposed an architecture review method. More recently, Breck et al. (2017) in their seminal work, examined the production readiness of machine learning systems, proposing a rubric to reduce technical debt and evaluate readiness. Parnin et al. (2023) documented the difficulties of building a production-ready copilots, by interviewing the team that built various copilots at Microsoft. In their influential blog post, Yan et al. ([n. d.]) shed light on some of the challenges that impact the production-readiness of FMware. Recently, Nahar et al. (Nahar et al., 2024) investigated emerging challenges, particularly those related to quality assurance, faced by developers integrating foundation models (FMs) into existing software products. Their study, based on interviews with 26 developers and a survey of 332 participants at Microsoft, provides recommendations for the development and integration of FMs into software products. Unlike their work, which is limited to insights from practitioners within a single company, our study analyzes grey literature from diverse sources to identify recurrent issues affecting all stages of the FMware lifecycle.To the best of our knowledge, this is the first study to systematically identify and organize the challenges in building production-ready FMware, highlighting core challenges and offering a technology roadmap for researchers and practitioners.
2.2. Surveys in FMware research
While several surveys have examined FMware and the role of LLMs in software engineering, they predominantly focus on academic publications or practitioner interviews. Jin et al. (Jin et al., 2024) provide a comprehensive exploration of LLMs and their application as autonomous agents in software engineering. Similarly, Fan et al. (Fan et al., 2023) survey LLMs’ potential in software development activities, including debugging and refactoring, emphasizing hybrid approaches that combine classical software engineering techniques with LLMs.
Zhang et al. (Zhang et al., 2023) investigate the impact of LLMs across five phases of the software engineering lifecycle, leveraging academic datasets to highlight technical challenges such as model tuning and evaluation. Hou et al. (Hou et al., 2024) curated a systematic review of 395 research papers to understand how LLMs are used in various software engineering tasks and on which tasks LLMs have shown success. Jiang et al. (Jiang et al., 2024) focus on LLMs for code generation, presenting a taxonomy of existing approaches based on structured academic benchmarks such as HumanEval and MBPP.
Wang et al. (Wang et al., 2024a) and others emphasize the use of LLMs in software testing, particularly for unit test generation and debugging, using standard datasets like Defects4J. However, all of these studies share a common limitation: an overreliance on academic datasets and benchmarks without addressing the contributions of open-source projects, forums, and informal knowledge exchange. Different from the aforementioned studies, Among the existing studies, Chen et al. (Chen et al., 2025) provide the most relevant foundation for our work by focusing on the challenges faced by LLM application developers. They present a comprehensive empirical study, mining 29,057 posts from the OpenAI developer forum to construct a detailed taxonomy of challenges. Their taxonomy highlights critical issues such as prompt engineering, API limitations, rate constraints, and hallucination management. To validate their findings, they extend their analysis to GitHub issues for LLaMa and Gemini, demonstrating the generalizability of their taxonomy across platforms. While their study provides actionable insights for API-driven LLM development, it overlooks challenges unique to users of smaller, self-hosted FMs, agentic development, and quality assurance activities, which are more common in enterprise environments prioritizing cost, quality, trustworthiness, and data privacy.
Unlike prior studies, which primarily rely on academic publications, this paper synthesizes insights from diverse data sources, including open-source project working group meeting notes, industry reports, grey literature (including academic publications) and informal knowledge exchanges. This holistic approach allows us to uncover the practical challenges and strategies employed in real-world FMware development. To the best of our knowledge, this is the first study to integrate academic insights, community-driven innovations, and practitioner expertise to provide a comprehensive overview of the recurrent issues that hinder the productionization of FMware across its various lifecycle stages and present a technology roadmap to address them.
3. Methodology
We conducted a semi-structured thematic synthesis, inspired by thematic analysis (Butler et al., 2024; Cruzes and Dybå, 2010), to identify and organize key challenges in developing production-ready FMware. Unlike a full thematic analysis, which involves detailed iterative coding and theme validation, we focused on systematically collecting and analyzing relevant sources, identifying issues, and grouping them into themes through expert-driven discussions.
The decision to employ a semi-structured thematic synthesis was driven by several factors. First, the sheer volume and rapid growth of publications (a recent study by Fan et al. (Fan et al., 2024) finds more than 5,000 academic publications in FMware between 2017 and 2023) in the FMware domain and the sheer scope of data (including meeting notes from several open source projects and conference reports from over 3 years) made comprehensive collection and iterative coding impractical, especially given the time constraints faced by the authors as industry researchers. Second, we mitigated potential biases by triangulating insights from diverse sources, including industry discussions, conference reports, and academic papers, ensuring a balanced and comprehensive perspective. Third, this methodology emphasizes actionable insights and practical relevance, which are critical for addressing real-world challenges in productionizing FMware. While closely following key steps of thematic analysis, such as identifying and grouping issues into themes, we adapted the process by skipping exhaustive iterative coding and instead leveraging expert judgment to refine themes.
We acknowledge that this approach has limitations, such as the absence of formal coding and the potential for subjectivity. However, these trade-offs were necessary to address the vastness and complexity of the domain, allowing us to provide a structured and systematic exploration of key challenges in transitioning FMware from demos to production-ready systems.
@@ 3. Methodology
textbfStep 1: Data Collection. We collected data from various sources to gain insights into the challenges that researchers and practitioners encounter when productionizing FMware.
@@
  • OPEA Project Participation: As active participants, we played a significant role in the OPEA initiative (Open Platform for Enterprise AI (OPEA) (OPEA, [n. d.])). In particular, the last author leads OPEA’s research working group. OPEA is a global multi-company collaborative initiative focused on addressing the challenges of making FMware enterprise-ready. OPEA provides detailed frameworks, architectural blueprints, and a four-step assessment for evaluating FMware in terms of enterprise readiness. We used meeting notes (some of which are publicly available (atl, 2024)) from OPEA discussions, which involved 43 organizations (e.g., Intel, AMD, RedHat, Docket, JFrog, Anyscale, LlamaIndex and SAP).
@@
  • SPDX Working Group: First author co-led SPDX community meetings (an ISO/IEC 5692:2021 Software Bill of Materials standard (SPDX, 2021)) over the past three years, contributing to discussions on the AI Bill of Materials (AI BOM), which included discussing challenges in building production-ready FMware. The meeting notes, available publicly (SPDX, [n. d.]), were incorporated into our analysis.
@@
  • Inception and Launch of the AIware Conferences, summit and Bootcamp: Authors of this paper spearheaded the first ACM International Conference on AI-Powered Software (AIware) (aiw, 2024), co-located with FSE 2024, the FM+SE summit series (fms, 2024) and the AIware Bootcamp (Rajbahadur et al., 2024). Engaging with participants from Google, Microsoft, GitHub, and other industry and academic professionals provided valuable insights into the challenges of productionizing FMware, which we summarized in a report used in our analysis (Hassan et al., 2024a).
@@
  • Internal Industry Experience: Our collaborations with customers and development teams to understand FMware’s functional and non-functional requirements. Additionally, all authors contributed to developing and productionizing FMArts, a comprehensive lifecycle engineering platform for FMware (Hassan et al., 2024b). Our practical experience in creating FMware with FMArts and making it production-ready (Hassan et al., 2024b) and integrating it into Huawei products (e.g., Huawei Cloud) informed the thematic analysis.
  • In-depth Literature Review: We began with a review of grey literature, including blog posts, whitepapers, and developer forums, and noticed recurring discussions about the production readiness of FMware. Building on these insights, we conducted a focused review of academic papers to find supporting evidence. Through this activity, we provide a comprehensive analysis of the challenges in deploying FMware, grounding all issues in Section 4 with citations and minimizing subjective bias.
The collected materials were reviewed to identify recurring issues relevant to FMware production readiness. These insights formed the basis for subsequent grouping and theme development.
Step 2: Data Analysis and Grouping. After data collection, three of the four authors reviewed the materials to identify recurring issues. Instead of formal iterative coding, we relied on a collaborative and expert-driven approach to extract key challenges and categorize them. Issues were grouped into preliminary themes based on shared characteristics or relevance to specific stages of the FMware lifecycle (Figure 1). For example, issues related to Data and FM Alignment included challenges such as low data quality and insufficient data diversity. This approach emphasized practical categorization rather than exhaustive formal coding.
Step 3: Collaborative Discussion to Identify Themes. The authors collaboratively refined the initial groupings into overarching themes (which we present as challenges) through discussions informed by their domain expertise. These themes spanned multiple FMware lifecycle stages and were grounded in the collected data but did not follow the strict iterative refinement protocols of thematic analysis. Instead, the grouping process was guided by the practical relevance of the challenges and their applicability to real-world FMware development.
Step 4: Thematic Consolidation. In the final step, the authors revisited and refined the overarching themes to ensure they comprehensively represented the identified challenges. The process prioritized clarity and relevance for practitioners while acknowledging the absence of exhaustive formal coding. These themes are presented in Section 5, offering a systematic view of critical challenges in transitioning FMware from demos to production-ready systems.
4. Recurrent Issues in Productionizing FMware
@@
Figure 1. FMware Lifecycle.
4.1. FM(s) Selection
The first step in developing FMware is to select the FM(s) to be used. There exist over 600,000 open-source models available (among FMs and the more traditional machine learning models) (Santa, [n. d.]). Typically, the number of FMs to be selected depends on the complexity and functionality required by the FMware. For example, a simple chatbot may only need one FM, while complex systems might require multiple FMs for its different functionalities.
Difficulty balancing functional requirements with performance and costs. Selecting FMs for FMware requires balancing functionality with performance, infrastructure, and costs (Kamath et al., 2024). Demos favor large general-purpose models like ChatGPT-4, neglecting latency and resource limits due to the typically small number of required API calls. LLaMa-70B, for example, needs 140GB VRAM to be hosted locally (Chavan et al., 2024). In production, smaller FMs are preferred for cost, latency, and privacy reasons (especially in the context of edge devices, such as smartphones) (Li et al., 2023; Wan et al., 2023), though they may lack the full capabilities of larger FMs (Chen and Varoquaux, 2024; Irugalbandara et al., 2024). However, it is important to note that even hosting smaller models can also increase ownership costs (Kamath et al., 2024).
4.2. Data and FM Alignment
After selecting FMs, the next crucial step is data and model alignment. This process tailors the FM(s) to meet the specific requirements and goals of a given FMware application (Xia et al., 2024).
Data alignment curates specialized data through labeling, generation, debugging, and testing, especially for enterprise cases (Liu et al., 2024b; Albalak et al., 2024). Domain-specific datasets are created to ensure that an FMware can perform well on the use-case specific data. Model alignment adjusts FM parameters using techniques like supervised fine-tuning and Reinforcement Learning with Human Feedback (RLHF) to ensure the FMware behaves as intended (Shen et al., 2023).
Low data quality. High data quality is crucial for reliable FMware, as poor data leads to untrustworthy outputs (Albalak et al., 2024).